OpenStudy (anonymous):

What is the 20th term of the arithmetic sequence 3, -3, -9, ... ?

OpenStudy (anonymous):

a = 3 d = -3 -3 = -6 \[t_n = a + (n-1)d\] Put n = 20 here, \[\large t_{20} = 3 + (20-1)(-6)\] \[\large t_{20} = 3 - 114\] \[\large t_{20} = -111\]

OpenStudy (anonymous):

same thing :D but a bit different